1. two charged spheres q1 = 2.00x10^11 C; and q2=-5.00x^-8C are held fixed at positions 0.0500m apart. Calculate the magnitude of each electrostatic force between two spheres
  2. Determine the electric flux of an electric field E = 2000 N/C with an area of 10m^2 and 45 degree angle
  3. compute for the amount of charge (q) that passes through a conductor in 10 seconds with the amount of current equal to 20 a.
1
Expert's answer
2022-04-07T18:25:48-0400

1.

"F=\frac{kq_1q_2}{r^2}=3.6~\mu N,"

2.

"\Phi=EA\sin\alpha=14~kVm,"

3.

"q=It=200~C."